Circular No. 8059

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003U IN NGC 6365A Further to IAUC 8057, T. Boles reports the discovery of an apparent

supernova (mag 18.0) on an unfiltered CCD image taken on Jan. 27.153 and 28.069 UT. The new object is located at α = 17^{h}22^{m}45^{s}.66, δ = +62^{o}09′50′′.4, which is ≈ 12′′.7 east and 7′′.3 south of the center of NGC 6365A. SN 2003U is not present on Boles’ images from 2002 Sept. 10 (limiting mag 19.0), Aug. 15, or July 26, and it is not present on Palomar Sky Survey red (1993) and blue (1989) plates.

SUPERNOVA 2003M IN UGC 7224 Further to IAUC 8058, T. Matheson et al. add that a spectrum of

SN 2003M, obtained by M. Calkins on Jan. 26.38 UT, shows it to be an unusual type-Ia supernova (not type Ic, as suggested on IAUC 8057): “As UGC 7224 is an elliptical galaxy, this would be an extraordinary event if it were a core-collapse supernova. The spectrum is like that of SN 2002jm (cf. IAUC 8031), reported by Foley et al. as similar to SN 1991bg at 3 weeks past maximum, but without the Na I D absorption seen in SN 1991bg.”

SUPERNOVA 2002gz M. Hamuy and M. Phillips, Carnegie Observatories; and J. Maza, Uni-

versity of Chile, report spectroscopic observations (range 320–920 nm) made of SN 2002gz between 2002 Oct. 29 and 2003 Jan. 10 UT with the Las Campanas Observatory Baade 6.5-m and Dupont 2.5-m telescopes: “A spectrum obtained on Nov. 10 bears some resemblance to that of the type-Ic supernova 1994I taken 4 days before maximum (Filippenko 1997, ARAA 35, 309), though a few differences are evident, such as (1) SN 2002gz was much bluer and showed no evidence for interstellar Na I D 589.3-nm absorption;

(2) the Ca H and K lines in SN 2002gz were much weaker; and (3) the BVI light curves of SN 2002gz appear to have been already declining by this date (i.e., our spectrum most likely was obtained after maximum). By Jan. 10, SN 2002gz began to show Na I D absorption, and the spectrum was similar to that of SN 1994I taken 26 days after maximum, but the Ca H and K lines in SN 2002gz were weak. The first spectrum of SN 2002gz obtained on Oct. 29 has a very blue continuum, broad low-contrast undulations, and a weak emission at the approximate position expected for Hα — perhaps suggesting an intermediate IIc type. An intermediate IIb type, as suggested on IAUC 8014 seems unlikely, since none of the spectra of SN 2002gz show evidence for strong He I lines.”

2003 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green